Molly has $15.05 in coins.
She has three more nickels than dimes.
She has six fewer quarters than dimes.
How many quarters does she have?

Answer by macston(5194)

Q=number of quarters D=no of dimes=Q+6; N=no of nickels=D+3=Q+9
$0.25Q+$0.10D+$0.05N=$15.05 Substitute for D and N
$0.25Q+$0.10(Q+6)+$0.05(Q+9)=$15.05
$0.25Q+$0.10Q+$0.60+$0.05Q+$0.45=$15.05
$0.40Q+$1.05=$15.05 Subtract $1.05 from each side
$0.40Q=$14.00 Divide each side by $0.40
Q=35 ANSWER She has 35 quarters.
CHECK
D=Q+6=35+6=41 She has 41 dimes
N=Q+9=35+9=44 She has 44 nickels
$25Q+$0.10D+$0.05N=$15.05
$0.25(35)+$0.10(41)+$0.05(44)=$15.05
$8.75+$4.10+$2.20=$15.05
$15.05=$15.05